Calculate a patient's intake for 4 hours if the patient receives-50 ml/hr by intravenous; 1 8 ounce glass of water; 2 6 ounce cups of coffee. How many milliters did the patient receive? Thanks.....Please show your work. Lean
Answer by nerdybill(7384)

Calculate a patient's intake for 4 hours if the patient receives-50 ml/hr by intravenous; 1 8 ounce glass of water; 2 6 ounce cups of coffee. How many milliliters did the patient receive?
.
Consider:"the patient receives-50 ml/hr by intravenous" for 4 hours:
4 * 50 = 200ml
.
Now since:
1 US fluid ounce = 29.5735296 ml
.
Consider: the 8 ounces of water consumed (convert to ml)
8 * 29.5735296 = 236.5882368 ml
.
Consider the coffee (convert to ml):
2*6*29.5735296 = 354.8823552 ml
.
Adding them all together:
200 + 236.5882368 + 354.8823552 = 791.47 ml